# SINKT: A Structure-Aware Inductive Knowledge Tracing Model with Large Language Model

### v1

@@ Abstract.
ccs: Social and professional topics Student assessmentccs: Applied computing E-learning
@@ 2.1. Knowledge Tracing
Concurrently, attention has been paid to more nuanced behavioral data and advanced neural architectures. For instance, LBKT (Xu et al., 2023) analyzes a range of learning behaviors, containing the speed of responses, the number of attempts, and the use of hints. SAKT (Pandey and Karypis, 2019) and AKT (Ghosh et al., 2020) introduce attention mechanisms (Bahdanau et al., 2016) to focus on the most relevant aspects of a student’s past interactions. Additionally, innovative approaches such as CMKT (Zhang et al., 2023b)assess students’ dynamic mastery of concepts. MF-DAKT (Zhang et al., 2021), which incorporates multiple student-related factors with a dual attention mechanism, showcases the field’s advancement towards more granular and complex models. SFKT (Zhang et al., 2023a) also demonstrates the application of KT models to handle sequences of variable lengths. Alongside these developments, graph-based KT models like GKT (Nakagawa et al., 2019) and GIKT (Yang et al., 2020) employ GNN (Kipf and Welling, 2017) to leverage the relational data among concepts. These models map the complex network of knowledge, allowing predictions that consider not just the individual learning trajectory but also the intricate structure of the knowledge itself.
@@ 2.2. LLM-Enhanced User Modeling
Existing LLM-enhanced user modeling methods are mostly applied in the domain of recommender systems, which significantly from the KT task. In recommender systems, users’ interests typically remain stable over short periods, whereas in ITS, the knowledge states of students evolve dynamically as they engage with educational contents such as exercises. Additionally, there is strong structural information, such as interdependency and correlation among the concepts in ITS. Therefore, there is a need for specially designed models that can effectively track and adapt to the unique dynamics of students’ learning processes in ITS.
@@ 4.2. Structural Information Encoder
In the graph $\mathcal{G}$, denote neighbor questions of the concept $c_{i}$ as $\mathcal{N}^{c}_{c_{i}}$ and $\mathcal{N}^{q}_{c_{i}}$, respectively. The set of neighbor concepts of the question $q_{i}$ is denoted as $\mathcal{N}^{c}_{q_{i}}$. We apply three different Graph Attention Networks (GAT) to aggregate neighborhood information for vertexes. Concept-Concept GAT layer fuse neighbor concept representations of the target concept $c_{i}$. Specifically, the Concept-Concept GAT can be expressed as:
@@ 5.2. Baselines
  • DKT (Piech et al., 2015) is the first model that introduces deep learning methods into the KT task, which uses RNN to model students’ learning history.
  • DHKT (Wang et al., 2019) assumes that skill is an abstraction of question and thus leverages the hierarchical structure between questions and concepts based on DKT.
  • DKVMN (Zhang et al., 2017) adds key-value memory network on DKT. The key matrix is static to represent the relations between different skills, and the value matrix is dynamic to store students’ mastery.
  • SKVMN (Abdelrahman and Wang, 2019) adds Hop-LSTM on DKVMN for better sequential modeling.
@@
  • SAKT (Pandey and Karypis, 2019) is the first model to add an attention mechanism into KT models, which uses a scaled dot-product attention mechanism to learn the past interactions’ importance in predicting students’ current questions.
  • EERNNA (Su et al., 2018) uses the hidden states of RNN to represent students’ knowledge states and considers the different impacts of previous states on the current response prediction.
@@
  • AKT (Ghosh et al., 2020) employs a self-attention mechanism to model the forgetting effect of students’ memory.
@@ 5.3. Implementation Details
In our experiment, we choose the most recent 200 records of students. We choose BERT to be the textual encoder for three public datasets with concept text and choose Vicuna to be the textual encoder for the Programming dataset with concept and question text. We will compare the performance of different textual encoders in Section 5.4. The layer number of the graph encoder is chosen from $\{1,2,3\}$. The parameter $d$ is set to $256$. The learning rate is chosen from $\{0.0001,0.00005\}$ with a decay at each epoch. The optimizer is Adam (Kingma and Ba, 2014). For a fair comparison, the hyper-parameters of baselines are carefully chosen to have the best performance. Our model is implemented on PyTorch and will be publicly available upon the acceptance of this work.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679760ccs: Social and professional topics Student assessmentccs: Applied computing E-learning
@@ 2.1. Knowledge Tracing
Recently, attention has been paid to more nuanced behavioral data and advanced neural architectures. For instance, LBKT (Xu et al., 2023) analyzes a range of learning behaviors, containing the speed of responses, the number of attempts, and the use of hints. SAKT (Pandey and Karypis, 2019) and AKT (Ghosh et al., 2020) introduce attention mechanisms (Bahdanau et al., 2016) to focus on the most relevant aspects of a student’s past interactions. Additionally, innovative approaches such as CMKT (Zhang et al., 2023)assess students’ dynamic mastery of concepts. MF-DAKT (Zhang et al., 2021), which incorporates multiple student-related factors with a dual attention mechanism, showcases the field’s advancement towards more granular and complex models. Alongside these developments, graph-based KT models like GKT (Nakagawa et al., 2019) and GIKT (Yang et al., 2020) employ GNN (Kipf and Welling, 2017) to leverage the relational data among concepts.
@@ 2.2. LLM-Enhanced User Modeling
Existing LLM-enhanced user modeling methods are mostly applied in the domain of recommender systems (RS), which differ significantly from the KT task. In RS, users’ interests typically remain stable over short periods, whereas in ITS, the knowledge states of students evolve dynamically as they engage with educational contents such as exercises. Additionally, there is strong structural information, such as interdependency and correlation among the concepts in ITS. Therefore, there is a need for specially designed models that can effectively track and adapt to the unique dynamics of students’ learning processes in ITS.
@@ 4.2. Structural Information Encoder
In the graph $\mathcal{G}$, denote neighbor questions of the concept $c_{i}$ be $\mathcal{N}^{c}_{c_{i}}$ and $\mathcal{N}^{q}_{c_{i}}$, respectively. The set of neighbor concepts of the question $q_{i}$ is denoted as $\mathcal{N}^{c}_{q_{i}}$. We apply three different Graph Attention Networks (GAT) to aggregate neighborhood information for vertexes. Concept-Question GAT layer fuse neighbor concept representations of the target concept $c_{i}$. Specifically, the Concept-Concept GAT can be expressed as:
@@ 5.2. Baselines
  • DKT (Piech et al., 2015) and DHKT (Wang et al., 2019) use deep learning methods on the KT task, which uses RNN to model students’ learning history.
  • DKVMN (Zhang et al., 2017), SKVMN (Abdelrahman and Wang, 2019) apply key-value memory network on the KT task.
@@
  • SAKT (Pandey and Karypis, 2019), EERNNA (Su et al., 2018) and AKT (Ghosh et al., 2020) use attention mechanism to learn the past interactions’ importance in predicting students’ current questions.
@@ 5.3. Implementation Details
In our experiment, we choose the most recent 200 records of students. We choose BERT to be the textual encoder for three public datasets with concept text and choose Vicuna to be the textual encoder for the Programming dataset with concept and question text. We will compare the performance of different textual encoders in Section 5.4. The layer number of the graph encoder is chosen from $\{1,2,3\}$. The parameter $d$ is set to $256$. The learning rate is chosen from $\{0.0001,0.00005\}$ with a decay at each epoch. The optimizer is Adam (Kingma and Ba, 2014). For a fair comparison, the hyper-parameters of baselines are carefully chosen to have the best performance. Our code is available now [^8] The PyTorch and MindSpore implementation is available at: https://github.com/tubehao/SINKT.git and https://github.com/mindspore-lab/models/tree/master/research/huawei-noah/SINKT..
@@ 6. Conclusion
Acknowledgements.
The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62177033, 62076161) and Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102). The work is also sponsored by Huawei Innovation Research Program. We thank MindSpore (min, 2020) for the partial support of this work.